# latexOverwriting file

### v1

@@ 4 Ensemble Kalman Filter
$v_{n+1}=\widehat{v}_{n+1}+\mathcal{C}^{vy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{\rm EK}_{n+1}\bigr{)}\mathcal{C}^{yy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{\rm EK}_{n+1}\bigr{)}^{-1}\bigl{(}{y}^{\dagger}_{n+1}-\widehat{y}_{n+1}\bigr{)},$
$(\widehat{v}_{n+1},\widehat{y}_{n+1})\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{\rm EK}_{n+1}.$
@@
The covariances $\mathcal{C}^{vy}$ and $\mathcal{C}^{yy}$ can be calculated exactly from the covariance of $\widehat{v}_{n+1}$ with $h(\widehat{v}_{n+1})$, from the covariance of $h(\widehat{v}_{n+1})$ with itself and from $\Gamma$. The formulae are the $J\to\infty$ limit of those given in (3).
Note that the map $v_{n}\mapsto v_{n+1}$ is nonlinear, stochastic and nonautonomous in the sense that it depends on the observed data. But, furthermore, the map depends on the law of $v_{n}$ since knowledge of this law is required to define ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{\rm EK}_{n+1}.$ Thus we now focus on finding an evolution equation for the law of $v_{n}$, which we denote by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n}.$
@@
$\mathfrak{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}},{y}^{\dagger})$ $\colon\mathbb{R}^{d}\times\mathbb{R}^{K}\to\mathbb{R}^{d};$
@@
$(v,y)$ $\mapsto v+\mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}})\mathcal{C}^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}})^{-1}\bigl{(}{y}^{\dagger}-y\bigr{)},$
@@
noting that this map is linear for any given pair $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}},{y}^{\dagger})\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})\times\mathbb{R}^{K}$. Recalling the notation $F_{\sharp}{\nu}$ for pushforward of a probability measure $\nu$ under a map $F$, we define
@@
$\mathsf{T}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}};{y}^{\dagger})$ $=\bigl{(}\mathfrak{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}},{y}^{\dagger})\bigr{)}_{\sharp}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}}.$
@@
Notice that $\mathsf{T}$, like $\mathsf{B}$, is nonlinear as a map on probability measures. With the definition of $\mathsf{T}$ it then follows [2] that
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n+1}=\mathsf{T}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n};{y}^{\dagger}_{n+1}),\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{0}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{0}.$
@@
The mean-field map $v_{n}\mapsto v_{n+1}$ is now well-defined by coupling it to this nonlinear evolution equation for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n}$, the law of $v_{n}.$ We refer to it as a mean-field map precisely because of this dependence on its own law.
@@
The key question we wish to address is the relationship between ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$. To focus on this question we write their evolution equations in parallel:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n+1}$ $=\mathsf{T}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n};{y}^{\dagger}_{n+1}),$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{B}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n};{y}^{\dagger}_{n+1}).$
@@
From this it is clear that the key is to understand when $\mathsf{T}\approx\mathsf{B}.$ In fact $\mathsf{T}\equiv\mathsf{B}$ on the set of Gaussian measures $\mathcal{G}(\mathbb{R}^{d}\times\mathbb{R}^{K})$ [2]. This is related to the fact that, in the Gaussian case, the map $\mathsf{T}$ effects an exact transport from the joint distribution of state and data to the conditional distribution of state given data. To obtain an error bound between ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ it is helpful to define
@@
$\mathsf{G}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}$ $=\operatorname*{arg\,min}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{p}}\in\mathcal{G}}\,\,d_{\rm KL}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{p}}).$
It then follows that $\mathsf{G}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}=\mathcal{N}({\rm mean}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}},{\rm cov}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}})$ [19, Theorem 4.7].
@@
The map $\mathsf{T}$ is not an optimal transport, in the sense of transporting one measure into another at minimal cost [22], but it is a transport that is well-adapted to numerical implementation. Links between transport and data assimilation were pioneered by Sebastian Reich; see [18] and the references therein.
@@
$d_{g}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\nu}})=\sup_{|f|\leqslant g}\bigl{\lvert}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}[f]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\nu}}[f]\bigr{\rvert},$
with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}[f]$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\nu}}[f]$ defined as before. In order to state our theorem about closeness of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n}$ we introduce the following measure of how close the true filter $\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}\}$ is to being Gaussian, in the lifted space of state and data:
@@
$\varepsilon:={\rm sup}\,_{0\leqslant n\leqslant N}\,\,d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n},\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}).$
For precise statement of the following theorem, and details of the conditions under which it holds, see [3].
@@
Let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{0}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{0}$ and assume that $\|\Psi\|_{L^{\infty}},\|h\|_{L^{\infty}}$ and $|h|_{C^{0,1}}$ are bounded by $r$. Then there is $C:=C(N,r)\in(0,\infty)$ such that
$\sup_{0\leqslant n\leqslant N}d_{g}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm EK}_{n})\leqslant C\varepsilon.$
@@
  • Consistency: in [3] a class of problems is identified within which $\varepsilon\to 0$ showing that in this setting the mean-field ensemble Kalman filter accurately approximates the true filter; at $\varepsilon=0$ a Gaussian problem is obtained and so the result concerns a class of problems in a small neighbourhood of Gaussians.

### v3

@@ 4 Ensemble Kalman Filter
$v_{n+1}=\widehat{v}_{n+1}+\mathcal{C}^{vy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{\rm EK}_{n+1}\bigr{)}\mathcal{C}^{yy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{\rm EK}_{n+1}\bigr{)}^{-1}\bigl{(}{y}^{\dagger}_{n+1}-\widehat{y}_{n+1}\bigr{)},$
$(\widehat{v}_{n+1},\widehat{y}_{n+1})\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{\rm EK}_{n+1}.$
@@
The covariances $\mathcal{C}^{vy}$ and $\mathcal{C}^{yy}$ can be calculated exactly from the covariance of $\widehat{v}_{n+1}$ with $h(\widehat{v}_{n+1})$, from the covariance of $h(\widehat{v}_{n+1})$ with itself and from $\Gamma$. The formulae are the $J\to\infty$ limit of those given in (4.1).
Note that the map $v_{n}\mapsto v_{n+1}$ is nonlinear, stochastic and nonautonomous in the sense that it depends on the observed data. But, furthermore, the map depends on the law of $v_{n}$ since knowledge of this law is required to define ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{\rm EK}_{n+1}.$ Thus we now focus on finding an evolution equation for the law of $v_{n}$, which we denote by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n}.$
@@
$\mathfrak{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}},{y}^{\dagger})$ $\colon\mathbb{R}^{d}\times\mathbb{R}^{K}\to\mathbb{R}^{d};$
@@
$(v,y)$ $\mapsto v+\mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}})\mathcal{C}^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}})^{-1}\bigl{(}{y}^{\dagger}-y\bigr{)},$
@@
noting that this map is linear for any given pair $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}},{y}^{\dagger})\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})\times\mathbb{R}^{K}$. Recalling the notation $F_{\sharp}{\nu}$ for pushforward of a probability measure $\nu$ under a map $F$, we define
@@
$\mathsf{T}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}};{y}^{\dagger})$ $=\bigl{(}\mathfrak{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}},{y}^{\dagger})\bigr{)}_{\sharp}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}}.$
@@
Notice that $\mathsf{T}$, like $\mathsf{B}$, is nonlinear as a map on probability measures. With the definition of $\mathsf{T}$ it then follows [2022arXiv220911371C] that
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n+1}=\mathsf{T}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n};{y}^{\dagger}_{n+1}),\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{0}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{0}.$
@@
The mean-field map $v_{n}\mapsto v_{n+1}$ is now well-defined by coupling it to this nonlinear evolution equation for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n}$, the law of $v_{n}.$ We refer to it as a mean-field map precisely because of this dependence on its own law.
@@
The key question we wish to address is the relationship between ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$. To focus on this question we write their evolution equations in parallel:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n+1}$ $=\mathsf{T}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n};{y}^{\dagger}_{n+1}),$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{B}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n};{y}^{\dagger}_{n+1}).$
@@
From this it is clear that the key is to understand when $\mathsf{T}\approx\mathsf{B}.$ In fact $\mathsf{T}\equiv\mathsf{B}$ on the set of Gaussian measures $\mathcal{G}(\mathbb{R}^{d}\times\mathbb{R}^{K})$ [2022arXiv220911371C]. This is related to the fact that, in the Gaussian case, the map $\mathsf{T}$ effects an exact transport from the joint distribution of state and data to the conditional distribution of state given data. To obtain an error bound between ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ it is helpful to define
@@
$\mathsf{G}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}$ $=\operatorname*{arg\,min}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{p}}\in\mathcal{G}}\,\,d_{\rm KL}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{p}}).$
It then follows that $\mathsf{G}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}=\mathcal{N}({\rm mean}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}},{\rm cov}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}})$ [2018arXiv181006191S, Theorem 4.7].
@@
The map $\mathsf{T}$ is not an optimal transport, in the sense of transporting one measure into another at minimal cost [MR2459454], but it is a transport that is well-adapted to numerical implementation. Links between transport and data assimilation were pioneered by Sebastian Reich; see [reich2015probabilistic] and the references therein.
@@
$d_{g}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\nu}})=\sup_{|f|\leqslant g}\bigl{\lvert}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}[f]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\nu}}[f]\bigr{\rvert},$
with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}[f]$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\nu}}[f]$ defined as before. In order to state our theorem about closeness of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n}$ we introduce the following measure of how close the true filter $\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}\}$ is to being Gaussian, in the lifted space of state and data:
@@
$\varepsilon:={\rm sup}\,_{0\leqslant n\leqslant N}\,\,d_{g}(\mathsf{G}\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n},\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}).$
For precise statement and proof of the following theorem, and in particular details of the conditions under which it holds, see [carrillo2022ensemble].
@@
Let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{0}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{0}$ and assume that $\|\Psi\|_{L^{\infty}},\|h\|_{L^{\infty}}$ and $|h|_{C^{0,1}}$ are bounded by $r$. Then there is $C:=C(N,r)\in(0,\infty)$ such that
$\sup_{0\leqslant n\leqslant N}d_{g}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm EK}_{n})\leqslant C\varepsilon.$
@@
  • Consistency: in [carrillo2022ensemble] a class of problems is identified within which there are sequences of problems along which $\varepsilon\to 0$. This demonstrates that there are problem classes within which the mean-field ensemble Kalman filter accurately approximates the true filter; at $\varepsilon=0$ a Gaussian problem is obtained and so the result concerns a class of problems in a small neighbourhood of Gaussians.